OpenStudy (naveenbhatia1312):

Find the exact value by using a half-angle identity. (5 points) cos 5pi/12

OpenStudy (mathmale):

Type out the "half-angle identity" for the cosine function, please.

OpenStudy (mathmale):

Think: \[\frac{ 5\pi }{ 12 }\]

OpenStudy (mathmale):

...is half of what angle?

OpenStudy (naveenbhatia1312):

10Pi

zepdrix (zepdrix):

5pi/12 is half as large as 5pi/6 So we can do something clever :) \[\large\rm \cos\left(\frac{5\pi}{12}\right)=\cos\left(\frac{\color{orangered}{5\pi/6}}{2}\right)\]

zepdrix (zepdrix):

And then you remember your Half-Angle Identity, yes?\[\large\rm \cos\left(\frac{\color{orangered}{x}}{2}\right)=\sqrt{\frac{1+\cos\color{orangered}{x}}{2}}\]

OpenStudy (naveenbhatia1312):

yes

OpenStudy (naveenbhatia1312):

i got to there but I can't come up with an answer

OpenStudy (naveenbhatia1312):

@zepdrix